OpenStudy (anonymous):

Which of the following equations is equivalent to 1/4x - 1/2y = 8? 2x - 4y = 8 2x - 4y = 64 4x - 2y = 64

OpenStudy (anonymous):

Can you help or not?

OpenStudy (daisyduck04):

oh sorry

OpenStudy (daisyduck04):

thats why you asked

OpenStudy (anonymous):

Yess

OpenStudy (daisyduck04):

no i was asking why like how did u find me i was on my smartscore and thats how i saw the message but i went back to this tab then i saw that i was on your question so yeah

OpenStudy (anonymous):

I multiplied the entire equation by 2 which resulted in x -2y = 32

OpenStudy (anonymous):

Ohhhh okay. I just happened to find you tbh.

OpenStudy (daisyduck04):

32*2 is

OpenStudy (anonymous):

64 lol

OpenStudy (anonymous):

But why multiply by 2 again?

OpenStudy (anonymous):

2x - 4y = 8 2x - 4y = 64 4x - 2y = 64 so it's the second one?

OpenStudy (daisyduck04):

ok so your answer is b

OpenStudy (anonymous):

Yeah, but I'm asking, why multiply the entire equation by 2 twice?

OpenStudy (anonymous):

Just so i can end up with one of the options?

OpenStudy (daisyduck04):

you multiplied

OpenStudy (daisyduck04):

bye sorry gtg

OpenStudy (anonymous):

Kay. But you didn't explain much..bye